I don't under Stan what to do in this problem
4/7+5/3

Answer by stanbon(75887)

4/7+5/3
----
You cannot add the fractions when they do not
have the same denominator:
---
least common denominator: 3*7 = 21
---
Convert each fraction to one which has 21 as its denominator:
-----
= (4/7)(3/3) + (5/3)(7/7)
-----
= 12/21 + 35/21
------
Combine the numerators over the lcd:
= (12+35)/21
---
= 47/21
=============
Cheers,
Stan H.
=============